OpenStudy (anonymous):

Evaluate the indefinite integral: ∫9dx= + C.

zepdrix (zepdrix):

The indefinite integral of a constant? Hmm.. What's the derivative of \(\large x\)? Just \(\large 1\) right? How bout the derivative of \(\large 9x\)? Hmm \(\large 9\) yes? So integrating \(\large 9\) should give us??

zepdrix (zepdrix):

If you're not quite convinced, you could also try applying the `Power Rule for Integration`. You could rewrite your integral like this,\[\large \int\limits 9x^0dx\]Then applying the power rule gives us,\[\large \frac{9x^{0+1}}{0+1}+C \qquad = \qquad 9x^1+C\]

OpenStudy (anonymous):

ok so i wasnt in class so how do we get this answer

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Integration is the opposite process of taking a derivative. Another word that you might here is "Anti-differentiation". \[\large \int\limits 9\;dx\] To evaluate this integral, we'll have to find the anti-derivative. When we integrate, the big swirly bar and the dx will disappear as a part of this process. \[\large \int\limits\limits 9\;dx \quad = \quad 9x\]See how the dx and swirly S disappeared? The derivative of \(\large 9x\) is \(\large 9\), therefore the anti-derivative of \(\large 9\) must be \(\large 9x\). Make a little bit of sense? D:

OpenStudy (anonymous):

ah ok ya that makes sense

zepdrix (zepdrix):

There is a +C is for another reason. Here is a quick example: \[\large f(x)=x^2+7\]\[\large f'(x)=2x\] Let's look at another similar function.\[\large g(x)=x^2+33\]\[\large g'(x)=2x\] So if I were to ask you to evaluate this integral,\[\large \int\limits 2x\;dx\]I'm asking you to find the anti-derivative of this function ~ The function that if we took the derivative of, would give us 2x. We have a problem though, is the anti-derivative going to be, \(\large x^2+7\), or \(\large x^2+33\) ??? The anti-derivative represents a FAMILY of solutions. It's the entire collection of possible x^2's. So we throw the +C on the end to show that the anti-derivative could be \(\large x^2+7\) or \(\large x^2+392\) or any unknown constant that would leave us with 2x if we took the function's derivative.

OpenStudy (anonymous):

ok that makes alot more sense

OpenStudy (anonymous):

thank you so much

zepdrix (zepdrix):

c: